OpenStudy (anonymous):

f(z)= z^2+1/3z Power and polynomial derivatives

OpenStudy (anonymous):

all iknw is that the derivative of \[f(z)=z^2+(1/3)z^{-1}\] is \[2z-(1/3)z^{-2}\] nt sure about power and polynomial der

OpenStudy (anonymous):

the answer in the back of the book is (z^2-1) /3z^2 Can you tell me the steps needed to get there I don't know the algebra :/

OpenStudy (anonymous):

well do you get the direvative answer i wrote on top

OpenStudy (anonymous):

is the question\[f(z)=z^2+\frac{ 1 }{ 3 z}\]

OpenStudy (anonymous):

i think it must be \[\frac{z^2+1}{3z}\]

OpenStudy (anonymous):

you have a choice, you can divide term by term and find the derivative my guess is that is what you are asked to do

OpenStudy (anonymous):

@satellite73 is answering the right question and youll get that answer on your txt book

OpenStudy (anonymous):

divide and get \[\frac{1}{3}z+\frac{1}{3z}\] derivative of \(\frac{1}{3}z\) is \(\frac{1}{3}\) derivative of \(\frac{1}{3z}\) is \(-\frac{1}{3z^2}\)

OpenStudy (anonymous):

algebra is \[\frac{1}{3}-\frac{1}{3z^2}=\frac{z^2}{3z^2}-\frac{1}{3z^2}=\frac{z^2-1}{3z^2}\]

OpenStudy (anonymous):

thanks Im reviewing all of this

OpenStudy (anonymous):

I guess I dont see how to get 1/3z + 1/3z